# VFEFL: Privacy-Preserving Federated Learning against Malicious Clients via Verifiable Functional Encryption

### v2

@@ VIII-B Experimental Results
As shown in Figures 7, 8, and 9, VFEFL consistently achieves high accuracy under LF attacks, demonstrating strong robustness. On the CIFAR10 dataset (Fig. 9), model convergence is relatively slower due to the aggregation rule normalizing each local model’s norm based on the baseline. This design mitigates malicious clients’ attempts to amplify local model parameters, as evidenced by the scaling attack results in Figures 4b, 5b, and 6b, but increases convergence time. Nevertheless, after approximately 20 iterations, our scheme attains satisfactory accuracy. Moreover, on MNIST and Fashion-MNIST, the global model maintains high classification accuracy and exhibits superior defense and stability against attacks compared to other methods. These results further confirm the robustness and effectiveness of our approach in adversarial settings.
@@ IX CONCLUSION
In this paper, we propose a novel federated learning framework, VFEFL, to address the limitation that existing frameworks depend on the assumption of two non-colluding servers or trusted third parties. Specifically, we propose a new aggregation rule and a decentralized verifiable functional encryption scheme, and rigorously prove their security. Furthermore, experiments on real datasets demonstrate that our scheme achieves robustness and fidelity comparable to existing methods, while eliminating reliance on such trust assumptions and providing strong verification guarantees.
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant 62372103, the Natural Science Foundation of Jiangsu Province under Grant BK20231149, the Jiangsu Provincial Scientific Research Center of Applied Mathematics under Grant BK20233002, and the Start-up Research Fund of Southeast University under Grant RF1028623300
@@ Acknowledgement
$\left|\Pr\left[\begin{array}[]{l}\begin{matrix}(\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p})\\ \leftarrow\texttt{Gen}(\lambda,p),\\ t\leftarrow\mathcal{D}_{p},\,h_{p}=\hat{h}_{p}^{t},\,x,x^{\prime}\\ \leftarrow D,\,b\overset{\$}{\leftarrow}\{0,1\},\\ Z_{0}=h^{x},\,Z_{1}=h^{x^{\prime}},\,b^{\prime}\leftarrow\\ \mathcal{A}(p,\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p},Z_{b},\texttt{Solve}(\cdot))\end{matrix}:b=b^{\prime}\end{array}\right]-\frac{1}{2}\right|.$
@@
From [14], one can set $S=2^{\lambda-2}\tilde{s}$, and instantiate $\mathcal{D}_{p}$ as the uniform distribution on $\{0,\cdots,S\}$ and $\mathcal{D}$ as the uniform distribution on $\{0,...,pS\}$.
@@
||| $\Pr\left[(g^{a},g^{b},g^{ab})\mid a,b\leftarrow\mathbb{Z}_{p}\right]$
@@
Definition 5 (Multi DDH Assumption[12])
Let $\mathcal{G}=(\mathbb{G},p,g)$, $g$ be a generator of $\mathbb{G}$ and $\mathrm{Adv}_{\mathbb{G}}^{ddh}(t)$ denote the best advantage of breaking the DDH assumption in $\mathbb{G}$ within time $t$. Given $X=g^{x},Y_{j}=g^{y_{j}}\overset{\$}{\leftarrow}\mathbb{G},j=1,\dots,m$, for any distinguisher $\mathcal{A}$ running within time $t$, the best advantage $A$ can get in distinguishing $\mathcal{D}_{m}=\{X,(Y_{j},g^{xy_{j}})_{j}\}$ from $\mathcal{D}^{\prime}_{m}=\{(X,(Y_{j},Z_{j})_{j})\mid Z_{j}\overset{\$}{\leftarrow}\mathbb{G},j=1,\dots,m\}$ is bounded by $\operatorname{Adv}^{\operatorname{ddh}}(t+4m\times t_{\mathbb{G}})$ (for a detailed proof, see [12]), where $t_{\mathbb{G}}$ is the time for an exponentiation in $\mathbb{G}$. Namely,
@@
  • Initialization: The challenger $\mathcal{C}$ runs algorithm SetUp($\lambda$) and KeyGen($pp$) to get public parameters $pp$ and the keys $\left((sk_{i},ek_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$ and chooses a random bit $b\overset{\$}{\leftarrow}\{0,1\}$. It sends $\left(vk_{CT},vk_{DK},pk\right)$ to the $\mathcal{A}$. $\mathcal{A}$ submits the indexs $\mathcal{CC}$ of the corruption queries for QCorrupt.
@@
  • Functional decryption key queries QDKeyGen: $\mathcal{A}$ can access to QDKeyGen with submitting $\left(i,y_{i},\ell_{y}\right)$ up to $q_{K}$ times. $\mathcal{C}$ runs $\left(dk_{y,i},\pi_{DKeyGenShare,i}\right)\leftarrow\texttt{DKeyGenShare}(sk_{i},pk,y_{i},\ell_{y},pp)$ and sends them to $\mathcal{A}$. $\mathcal{A}$ obtaining the decryption key validation key $vk_{DK}$ in the Initialization phase can implement the validation itself without querying VerifyDK($(dk_{i,y})_{i\in[n]},vk_{DK},pp$).
@@
  • Finalize: $\mathcal{A}$ outputs its guess $b^{\prime}$ on the bit $b$, and this process produces the outcome $\beta$ of the security game, based on the analysis provided below. Let $\mathcal{CC}$ represent the set of corrupted clients (the indices $i$ that are input to QCorrupt throughout the entire game), while $\mathcal{HC}$ denotes the set of honest (noncorrupted) clients. We set output $\beta=(b^{\prime}==b)$ if none of the following occur, otherwise, $\beta\overset{\$}{\leftarrow}\{0,1\}$
@@
  • Encryption queries QEncrypt: $\mathcal{A}$ has $q_{E}$ access to QEncrypt to obtain a correct encryption with submitting an index $i$, message $x_{i}\in\mathcal{M}^{m}$ and labels $\ell$ and $\left(\ell_{Enc,j}\right)_{j\in[m]}$. $\mathcal{C}$ runs Encrypt($ek_{i},x_{i},x_{0},\ell,\left(\ell_{Enc,j}\right)_{j\in[m]}$) to gain $\left(C_{\ell,i},\pi_{Encrypt,i}\right)$ and sends them to $\mathcal{A}$. $\mathcal{A}$ obtaining the validation key $vk_{CT}$ in the key generation phase can implement the ciphertext validation itself without querying VerifyCT($C_{\ell,i},vk_{DK}$).
  • Functional key share queries QDKeyGen: $\mathcal{A}$ has $q_{K}$ access to QDKeyGen to gain a correct functional key share with sunbmitting $\left(i,\ell_{y},y\right)$. $\mathcal{C}$ generates $\left(dk_{y,i},\pi_{DKeyGenShare,i}\right)$ by executing DKeyGenShare($sk_{i},pk,y_{i},\ell_{y},pp$) and sends them to $\mathcal{A}$. $\mathcal{A}$ obtaining the decryption key validation key $vk_{DK}$ in the key generation phase can implement the validation itself without querying VerifyDK($(dk_{i,y})_{i\in[n]},vk_{DK},pp$).
@@ Appendix B Instantiation of $\pi_{Encryption}$
$\rho\overset{\$}{\leftarrow}\mathbb{Z}^{2}_{p},\quad k\overset{\$}{\leftarrow}\mathbb{Z}_{p}^{m}$
@@
$\mathcal{V}:\alpha\overset{\$}{\leftarrow}\mathbb{Z}^{*}_{p}$
@@
$g^{t}(u^{m})^{\tau_{\alpha}}$ =?${=}T_{0}\left(T_{1}\right)^{\alpha}\left(T_{2}\right)^{\alpha^{2}}$
$g^{\langle L,x_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}$ =?${=}T_{0}^{y_{i}}(T^{\prime}_{1})^{\alpha}$
$w^{L}(u^{m})^{\omega}$ =?${=}V_{i}(K)^{\alpha}$
$t$ =?${=}\langle L,L\rangle$
$(v)^{\omega}$ =?${=}\mathrm{com}_{i}(\mathrm{com}^{*})^{\alpha}$
@@ Appendix C Instantiation of $\pi_{KeyGenShare}$
$r_{\hat{k},i},r_{s,i}\overset{\$}{\leftarrow}\mathbb{Z}_{p}^{2},\quad r_{t,i}\overset{\$}{\leftarrow}[0,2^{\lambda}pS]^{2}$
@@
$\mathcal{V}:\beta\overset{\$}{\leftarrow}\mathbb{Z}_{p}$
@@
$R_{T,i}$ =?${=}(T_{i,b}^{\beta}\cdot h_{p}^{z_{t,i,b}})_{b\in[2]}$
$K_{\Sigma,i}$ =${=}\left\{\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\right\}_{b\in[2]}\in G$
$R_{d,i}$ =?${=}\left(\mathrm{d}_{i,b}^{\beta}\cdot f^{z_{\hat{k},i,b}}K_{\Sigma,i,b}^{z_{t,i,b}}\right)_{b\in[2]}$
$R_{\mathrm{dk}_{i}}$ =?${=}\left((\mathrm{dk}_{i})^{\beta}\cdot(\hat{v}_{b})^{z_{\hat{k},i}}\cdot h^{z_{s,b}\cdot y_{i}}\right)_{b\in[2]}$
$R_{\mathrm{com},i}$ =?${=}(\mathrm{com}_{i})^{\beta}(v)^{z_{s,i}}$
@@
A formal proof are provided in [14].
@@ Appendix D Proof of Theorem 5
[…]

### v3

@@ VIII-B Experimental Results
As illustrated in Figures 7, 8 and 9, the proposed scheme consistently achieves high accuracy under LF attacks across all three datasets, while effectively suppressing the attack success rate. This highlights the robustness of our approach in defending against LF attacks. But in the CIFAR10 dataset, the global models obtained from the training of our scheme converge relatively slowly, as shown in Fig 9. This phenomenon is mainly due to the fact that the adopted aggregation rule normalises the norm of each local model based on the root model. This design effectively suppresses the malicious client’s strategy to enhance the attack effectiveness by amplifying the parameters of the local models, as demonstrated by the scaling attack in Figs 4b, 5b, and 6b. However, this normalisation mechanism also increases the number of iterations required for model convergence to some extent. Nevertheless, after about 20 iterations, the global model trained by our scheme is still able to achieve satisfactory accuracy. In addition, the global model trained by the proposed scheme not only maintains high classification accuracy on MNIST and Fashion-MNIST datasets, but also demonstrates better defence capability and performance stability against malicious attacks compared to other schemes in multiple iterations. The above results further validate the robustness and effectiveness of this paper’s method in malicious environments.
@@ IX CONCLUSION
In this paper, we presented VFEFL, based on decentralized verifiable functional encryption. To implement the framework, we proposed a new aggregation rule and a new verifiable functional encryption scheme, and proved their security. Our VFEFL scheme achieves privacy protection and resistance to Byzantine clients in a single-server setup without introducing additional trusted third parties, and is more applicable to general scenarios. We analyzed both the privacy and robustness of the scheme, in addition to the verifiability, fidelity and self-containment achieved by VFEFL. Finally, we further demonstrated the effectiveness of the VFEFL scheme through experiments.
@@
$\left|\Pr\left[\begin{array}[]{l}\begin{matrix}(\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p})\\ \leftarrow\texttt{Gen}(\lambda,p),\\ t\leftarrow\mathcal{D}_{p},\,h_{p}=\hat{h}_{p}^{t},\,x,x^{\prime}\\ \leftarrow D,\,b\overset{\mathdollar}{\leftarrow}\{0,1\},\\ Z_{0}=h^{x},\,Z_{1}=h^{x^{\prime}},\,b^{\prime}\leftarrow\\ \mathcal{A}(p,\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p},Z_{b},\texttt{Solve}(\cdot))\end{matrix}:b=b^{\prime}\end{array}\right]-\frac{1}{2}\right|.$
@@
From [VMCFE], one can set $S=2^{\lambda-2}\tilde{s}$, and instantiate $\mathcal{D}_{p}$ as the uniform distribution on $\{0,\cdots,S\}$ and $\mathcal{D}$ as the uniform distribution on $\{0,...,pS\}$.
@@
$|$ $\Pr\left[(g^{a},g^{b},g^{ab})\mid a,b\leftarrow\mathbb{Z}_{p}\right]$
@@
Definition 5 (Multi DDH Assumption[DMCFE])
Let $\mathcal{G}=(\mathbb{G},p,g)$, $g$ be a generator of $\mathbb{G}$ and $\mathrm{Adv}_{\mathbb{G}}^{ddh}(t)$ denote the best advantage of breaking the DDH assumption in $\mathbb{G}$ within time $t$. Given $X=g^{x},Y_{j}=g^{y_{j}}\overset{\mathdollar}{\leftarrow}\mathbb{G},j=1,\dots,m$, for any distinguisher $\mathcal{A}$ running within time $t$, the best advantage $A$ can get in distinguishing $\mathcal{D}_{m}=\{X,(Y_{j},g^{xy_{j}})_{j}\}$ from $\mathcal{D}^{\prime}_{m}=\{(X,(Y_{j},Z_{j})_{j})\mid Z_{j}\overset{\mathdollar}{\leftarrow}\mathbb{G},j=1,\dots,m\}$ is bounded by $\operatorname{Adv}^{\operatorname{ddh}}(t+4m\times t_{\mathbb{G}})$ (for a detailed proof, see [DMCFE]), where $t_{\mathbb{G}}$ is the time for an exponentiation in $\mathbb{G}$. Namely,
@@
  • Initialization: The challenger $\mathcal{C}$ runs algorithm SetUp($\lambda$) and KeyGen($pp$) to get public parameters $pp$ and the keys $\left((sk_{i},ek_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$ and chooses a random bit $b\overset{\mathdollar}{\leftarrow}\{0,1\}$. It sends $\left(vk_{CT},vk_{DK},pk\right)$ to the $\mathcal{A}$. $\mathcal{A}$ submits the indexs $\mathcal{CC}$ of the corruption queries for QCorrupt.
@@
[…]
@@
  • Finalize: $\mathcal{A}$ outputs its guess $b^{\prime}$ on the bit $b$, and this process produces the outcome $\beta$ of the security game, based on the analysis provided below. Let $\mathcal{CC}$ represent the set of corrupted clients (the indices $i$ that are input to QCorrupt throughout the entire game), while $\mathcal{HC}$ denotes the set of honest (noncorrupted) clients. We set output $\beta=(b^{\prime}==b)$ if none of the following occur, otherwise, $\beta\overset{\mathdollar}{\leftarrow}\{0,1\}$
@@
[…]
@@ Appendix B Instantiation of $\pi_{Encryption}$
$\rho\overset{\mathdollar}{\leftarrow}\mathbb{Z}^{2}_{p},\quad k\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{m}$
@@
$\mathcal{V}:\alpha\overset{\mathdollar}{\leftarrow}\mathbb{Z}^{*}_{p}$
@@
$g^{t}(u^{m})^{\tau_{\alpha}}$ $\overset{?}{=}T_{0}\left(T_{1}\right)^{\alpha}\left(T_{2}\right)^{\alpha^{2}}$
$g^{\langle L,x_{0}\rangle}(u^{m})^{\tau^{\prime}_{\alpha}}$ $\overset{?}{=}T_{0}^{y_{i}}(T^{\prime}_{1})^{\alpha}$
$w^{L}(u^{m})^{\omega}$ $\overset{?}{=}V_{i}(K)^{\alpha}$
$t$ $\overset{?}{=}\langle L,L\rangle$
$(v)^{\omega}$ $\overset{?}{=}\mathrm{com}_{i}(\mathrm{com}^{*})^{\alpha}$
@@ Appendix C Instantiation of $\pi_{KeyGenShare}$
$r_{\hat{k},i},r_{s,i}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2},\quad r_{t,i}\overset{\mathdollar}{\leftarrow}[0,2^{\lambda}pS]^{2}$
@@
$\mathcal{V}:\beta\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$
@@
$R_{T,i}$ $\overset{?}{=}(T_{i,b}^{\beta}\cdot h_{p}^{z_{t,i,b}})_{b\in[2]}$
$K_{\Sigma,i}$ ${=}\left\{\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\right\}_{b\in[2]}\in G$
$R_{d,i}$ $\overset{?}{=}\left(\mathrm{d}_{i,b}^{\beta}\cdot f^{z_{\hat{k},i,b}}K_{\Sigma,i,b}^{z_{t,i,b}}\right)_{b\in[2]}$
$R_{\mathrm{dk}_{i}}$ $\overset{?}{=}\left((\mathrm{dk}_{i})^{\beta}\cdot(\hat{v}_{b})^{z_{\hat{k},i}}\cdot h^{z_{s,b}\cdot y_{i}}\right)_{b\in[2]}$
$R_{\mathrm{com},i}$ $\overset{?}{=}(\mathrm{com}_{i})^{\beta}(v)^{z_{s,i}}$
@@
A formal proof are provided in [VMCFE].
@@ Appendix D Proof of Theorem 5
$L_{1}=x_{i}+k\alpha_{1}$ $\text{ and }L_{2}=x_{i}+k\alpha_{2}$
$\omega_{1}=s_{i}+\alpha_{1}\rho$ $\text{ and }\omega_{2}=s_{i}+\alpha_{2}\rho$
$\tau^{\prime}_{\alpha_{1}}=\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha_{1}$ $\text{ and }\tau^{\prime}_{\alpha_{2}}=\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha_{2}$
$\tau_{\alpha_{1}}=\tau_{0}+\tau_{1}\alpha_{1}+\tau_{2}\alpha_{1}^{2}$ $\text{ and }\tau_{\alpha_{2}}=\tau_{0}+\tau_{1}\alpha_{2}+\tau_{2}\alpha_{2}^{2},$